# Circumventing shortcuts in audio-visual deepfake detection datasets with unsupervised learning

### v1

@@ 4.2 Experimental setup
[…]
@@ 4.3 Experimental results
We train our method, AVH-Align, as well as its supervised variant, AVH-Align/sup, on the two datasets, FakeAVCeleb and AV-Deepfake1M, and evaluate on either of them (that is, both in-domain and out-of-domain). For each evaluation, in addition to the original set, we also consider the trimmed version of the validation dataset (trim: ). The trimmed variant is obtained by removing the starting 0.40 ms of audio and first video frame. We always use the untrimmed, original data for training. Results are shown in Table 2.
Impact of leading silence bias. We observe that AVH-Align is not impacted by removing the leading silence. In particular, AVH-Align shows even performance for FakeAVCeleb ($96.3\%$ AUC) and its trimmed version ($96.4\%$ AUC), as well as for AV-Deepfake1M ($85.9\%$ AUC) and its trimmed version ($86.2\%$ AUC). The supervised variant, AVH-Align/sup, shows a different behavior: on AV-Deepfake1M the performance drops by $15.7\%$ and $28.4\%$ in the trimmed case depending on the training dataset. On FakeAVCeleb removing the leading silence has a more negligible impact. Since FakeAVCeleb has only full manipulations, we hypothesize that there are other signals throughout the video that the classifier can pick up during learning, making it less prone to rely on the silence bias.
Comparison with other unsupervised methods. We compare our results with those of AVAD [16] which is also an usupervised method, trained on real data only. Similar to AVH-Align it is not impacted by the spurious leading silence, showing nearly identical results for trimmed and untrimmed datasets. However, AVAD’s overall performance is considerably worse on both datastes.
Score visualization. Figure 5 shows the per frame scores obtained by the two methods, AVH-Align and AVH-Align/sup, together with the groundtruth manipulated interval. For the AVH-Align method, the scores represent the misalignment probability between the audio and visual streams at each time frame; for the AVH-Align/sup method, the scores represent the probability of an audio-video frame of being fake. We see that AVH-Align/sup always predicts the first frame as being fake, confirming once again that it has learned to associate the spurious leading silence with a video being fake. When detected, the actual fake region is given a considerable lower fakeness probability than that assigned to the leading silence. On the other hand, AVH-Align is not affected by the presence of leading silence. In this case the manipulated areas have higher misalignment scores, a frame-level evaluation yielding 77.7% AUC. We do notice that there are other regions in the video that get a high misalignment score. This happens because the operations that introduce fake regions can introduce desynchronizations between the audio video streams even in areas that were not intentionally manipulated.
@@ 4.4 Comparison on the official AV-Deepfake1M test
Zero-shot methods
Video-LLaMA (13B) E5 [59] AV 50.74
CLAP [55] A 50.83
M2TR [52] V 50.18
LipForensics [17] V 51.57
Face X-Ray [27] V 61.54
SBI [47] V 55.10
Video-level methods
Meso4 [1] V 50.22
MesoInception [1] V 50.05
SBI [47] V 65.82
@@ 4.5 Further analysis
Ablation of AVH-Align components. Here we analyze the impact of various design choices on the performance of the AVH-Align method. Specifically, we investigate the following: i) removing feature normalization; ii) reducing the training set size from 45,000 to 8,782 samples (the real ones from the AVH-Align/sup train set); iii) using mean score pooling instead of log-sum-exp; iv) extracting features with the AV-HuBERT checkpoint pretrained on LRS3 only instead of LRS3 and VoxCeleb2; v) using a linear layer instead of MLP. Results are shown in Table 4. All ablations lead to a decrease in performance with the exception of mean score pooling for FakeAVCeleb. This is expected, however, since fake videos in FakeAVCeleb are manipulated at every frame, while AV-Deepfake1M has only locally-manipulated fake videos.
FakeAVCeleb AV-Deepfake1M
Configuration Trim: Trim: Trim: Trim:
AVH-Align 96.3 96.4 85.9 86.2
… Feature norm: 93.9 94.0 84.8 85.2
… Train size: 8.7k 90.5 91.2 82.5 82.5
… Score pooling: mean 98.3 98.3 76.8 76.8
… AV-HuBERT: LRS3 87.7 87.7 60.3 60.3
… Alignment net: Linear 34.3 33.8 51.5 51.4
Table 4: Ablation of the main components of AVH-Align. Mean score pooling is beneficial to FakeAvCeleb since it has full video manipulations. All other ablations degrade the performance.
Analysis of AVH-Align/sup architecture. We investigate how a simpler architecture impacts the supervised model. To this end, we train a linear layer on top of AV-HuBERT representations instead of the MLP network. We see in Table 5 that this modification has little impact on both FakeAVCeleb and AV-Deepfake1M. Unlike the unsupervised case, when a linear layer was too weak to learn, in the supervised case, a linear layer performs almost on par with the MLP. The largest difference between the MLP and the linear layer can be observed when training on FakeAVCeleb and testing on AV-Deepfake1M. With a linear layer, the performance drop from untrimmed to the trimmed AV-Deepafake1M is $21\%$ (from $85.1$ to $67.1$), while with a MLP, the drop is $28\%$ (from $88.8$ to $63.3$).
@@
Linear FakeAVCeleb 99.5 99.4 85.1 67.1
MLP FakeAVCeleb 99.5 99.4 88.8 63.6
Linear AV-Deepfake1M 76.2 71.2 99.9 83.3
MLP AV-Deepfake1M 75.9 72.2 99.9 84.2
Table 5: Architecture analysis of AVH-Align/sup. In the supervise case there is little impact of using a linear layer compared to a MLP.

### v3

@@ 4.2 Experimental setup
Datasets and metrics. We conducted our experiments on the two datasets introduced in Sec. 3.1: FakeAVCeleb and AV-Deepfake1M. For FakeAVCeleb, we use 70% of the dataset for training and validation, and the rest of 30% for testing. For AV-Deepfake1M, we evaluate on 10k samples from the original validation set, but for the best performing models we also report results on the official withheld test set. To train AVH-Align, we use 50k real samples from the training set of AV-Deepfake1M (45k samples for training and 5k for validation). Since AV-Deepfake1M is based on VoxCeleb2, these samples originate from VoxCeleb2. To train the supervised variant, AVH-Align/sup, on AV-Deepfake1M, we select the same number of samples (45k for training and 5k for validation), but this time coming from both the real and fake classes. For both datasets, we evaluate video-level detection. Following prior work, we report results in terms of the area under the receiver operator characteristic curve (AUC) and average precision (AP), with the fake class being the positive class.
@@ 4.3 Experimental results
We train our method, AVH-Align, as well as its supervised variant, AVH-Align/sup, on the two datasets, FakeAVCeleb and AV-Deepfake1M, and evaluate on either of them (that is, both in-domain and out-of-domain). For each evaluation, in addition to the original set, we also consider the trimmed version of the validation dataset (trim: ). The trimming is performed as described in Sec. 3.3. We always use the untrimmed, original data for training. Results are shown in Table 2.
Impact of leading silence bias. We observe that AVH-Align is robust to the leading silence bias: trimming the silence has no effect on FakeAVCeleb and only a slight effect on AV-Deepfake1M ($2.4\%$ AUC). On the other hand, the performance of the supervised variant, AVH-Align/sup, decreases considerably when testing on AV-Deepfake1M (by $16.9\%$ or $5.4\%$ AUC, depending on the train set). On FakeAVCeleb removing the leading silence has no impact for AVH-Align/sup, when training also on FakeAVCeleb. This may happen because FakeAVCeleb has full manipulations and there is other useful information in the video. Instead, when training on the locally-manipulated AV-Deepfake1M, the difference becomes significant ($6.7\%$ AUC).
Comparison with other unsupervised methods. We compare our results with those of AVAD [17] and SpeechForensics [30], which are also usupervised methods, trained on real data only. Similar to AVH-Align they are not impacted by the spurious leading silence, showing nearly identical results for trimmed and untrimmed datasets. However, their overall performance is considerably worse than ours on the AV-Deepfake1M dataset.
Score visualization. Figure 5 shows the per frame scores obtained by the two methods, AVH-Align and AVH-Align/sup, together with the groundtruth manipulated interval. For the AVH-Align method, the scores represent the misalignment probability between the audio and visual streams at each time frame; for the AVH-Align/sup method, the scores represent the probability of an audio-video frame of being fake. We see that AVH-Align/sup always predicts the first frame as being fake, confirming once again that it has learned to associate the spurious leading silence with a video being fake. When detected, the actual fake region is given a considerable lower fakeness probability than that assigned to the leading silence. On the other hand, AVH-Align is not affected by the presence of leading silence. In this case the manipulated areas have higher misalignment scores, a frame-level evaluation yielding 77.7% AUC. We do notice that there are other regions in the video that get a high misalignment score. This happens because the separation and re-composition of the audio and video streams can introduce desynchronizations between the two even in areas that were not intentionally manipulated.
@@ 4.5 Further analysis
Ablation of AVH-Align components. Here we analyze the impact of various design choices on the performance of the AVH-Align method. Specifically, we investigate the following: 1) removing feature normalization; 2) reducing the training set size from 45,000 to 8,782 samples (the real ones from the AVH-Align/sup train set); 3) using mean score pooling instead of log-sum-exp; 4) extracting features with the AV-HuBERT checkpoint pretrained on LRS3 only instead of LRS3 and VoxCeleb2; 5) using a linear layer instead of MLP. Results are shown in Table 4. All ablations lead to a decrease in performance with the exception of mean score pooling for FakeAVCeleb and AVLips. This is expected, however, since fake videos in FakeAVCeleb and AVLips are manipulated at every frame, while AV-Deepfake1M has only locally-manipulated fake videos.
Configuration FAVC AV1M AVLips
AVH-Align 94.6 85.9 86.3
1 … Feature norm: 92.4 84.8 86.1
2 … Train size: 8.7k 89.1 82.5 83.0
3 … Score pooling: mean 97.4 76.8 88.8
4 … AV-HuBERT: LRS3 84.8 60.3 80.1
5 … Alignment net: Linear 31.5 51.9 43.4
Table 4: Ablation of the main components of AVH-Align in terms of AUC. Mean score pooling helps when videos are fully generated (FAVC, AVLips). Other ablations degrade the performance.
Impact of VoxCeleb2 pretraining. Both FakeAVCeleb and AV1M datasets use VoxCeleb2 as their source dataset. Can there be a data leakage from AV-HuBERT VoxCeleb2 features to these datasets? We evaluate on a dataset that is not using VoxCeleb2, AVLips [31]. We observe that VoxCeleb2 features perform better even on this dataset, suggesting that these features are simply stronger than LRS3 features. This is may be explained by the size of VoxCeleb data (about four times larger than LRS3) and is in line with the results on other downstream tasks (e.g., lip reading [49]).
Analysis of AVH-Align/sup architecture. We investigate how a simpler alignment network impacts the supervised model. To this end, we train a linear layer on top of the AV-HuBERT representations instead of the MLP network. In Table 5, we observe that unlike the unsupervised case where a linear layer was too weak to learn, here a linear layer performs as well or even better than the MLP. The largest difference between the MLP and then linear layer is observed when training on FakeAVCeleb and testing on AV-Deepfake1M. With a linear layer, the performance drop from untrimmed to trimmed AV-Deepafake1M is $19.3\%$ (from $85.9\%$ to $66.6\%$ AUC), while with an MLP, the drop is $5.4\%$ (from $69.0\%$ to $63.6\%$ AUC).
@@
Linear FakeAVCeleb 99.6 99.5 85.9 66.6
MLP FakeAVCeleb 99.2 99.2 69.0 63.6
Linear AV-Deepfake1M 76.2 69.7 99.9 83.2
MLP AV-Deepfake1M 77.5 70.8 100.0 83.1
Table 5: Architecture analysis of AVH-Align/sup. In most cases, the linear and MLP architectures perform similarly.
@@ 6 Conclusions
Acknowledgments
This work was supported in part by the EU Horizon projects AI4TRUST (No. 101070190) and ELIAS (No. 101120237), and by CNCS-UEFISCDI (PN-IV-P7-7.1-PTE-2024-0600).